OpenStudy (anonymous):

What is the correct solution to the equation? 8(x + 2) = 24 A. 1 B. 3 C. 5 D. 64

OpenStudy (anonymous):

8(x + 2) = 24 8x + 16 = 24 8x = 24 - 16 8x = 8 x = 8/8 therefore x is 1 SO THE ANSWER IS A